OpenStudy (karatechopper):

can anyone give me standard form,slope intercept form, and point slope form questions? i want to study for my test. only basic ones no exponents or anythin like tht

hero (hero):

y = x^2-9 Find slope and intercept

OpenStudy (karatechopper):

i cant lern exponents!

OpenStudy (karatechopper):

u will confuse me!

hero (hero):

Wow, my whole picture showed up

hero (hero):

When did this feature get added?

OpenStudy (anonymous):

NOW

OpenStudy (karatechopper):

just 5 mins ago and can u give me a question tht i CAN do?!

hero (hero):

Nice. Now I can look at sweet lips without having to scroll up

OpenStudy (anonymous):

WHAT !?

OpenStudy (karatechopper):

sweet lips?!

hero (hero):

lol

hero (hero):

Meow, you're the one that posted a picture of your lips. Don't blame me for stating the obvious

OpenStudy (karatechopper):

i neeed a question not frickin sweet lips?!

OpenStudy (anonymous):

lol

OpenStudy (karatechopper):

ha hero

OpenStudy (anonymous):

LOL Heroine I'll Change it then =P

OpenStudy (karatechopper):

can i have a question now?! PLEASE! with a... aw nvrm the rest i need a question

hero (hero):

I just gave you one

OpenStudy (anonymous):

3y = 76x + 90

OpenStudy (karatechopper):

how do iwrite

OpenStudy (karatechopper):

what form?

OpenStudy (anonymous):

standard

hero (hero):

standard

OpenStudy (anonymous):

lol

OpenStudy (karatechopper):

k

OpenStudy (karatechopper):

76x-3y= -90

OpenStudy (anonymous):

ax + by + c = 0 is the standard form

hero (hero):

After you put it in standard form correctly, then put it in slope-intercept form

OpenStudy (anonymous):

Ok Done =)

hero (hero):

What happened to the sweet lips?

OpenStudy (karatechopper):

ishaan u r wrong

OpenStudy (karatechopper):

im correct

OpenStudy (karatechopper):

its Ax + By=C

OpenStudy (anonymous):

I covered it =D

hero (hero):

There are different ways of expressing standard form

OpenStudy (anonymous):

Hmm Agree with Hero

hero (hero):

Meow, do not cover your lips.

OpenStudy (karatechopper):

but i learned it one way!!!!

OpenStudy (anonymous):

aww =P

OpenStudy (karatechopper):

http://courses.wccnet.edu/~palay/precalc/22mt01.htm

OpenStudy (anonymous):

You're right

OpenStudy (karatechopper):

yes!!!

OpenStudy (karatechopper):

anymore>?

OpenStudy (anonymous):

You already Know it,Now What Karatechopper ?

OpenStudy (karatechopper):

o in slope intercept form like hero said

OpenStudy (karatechopper):

y=76/3x +30

OpenStudy (anonymous):

http://www.khanacademy.org/exercise/slope_of_a_line

OpenStudy (anonymous):

Khan Academy is Really helpfull

OpenStudy (karatechopper):

no i want u guys to give me some!

OpenStudy (anonymous):

Yeah it is Helpful Internet is full of Learning resources

OpenStudy (karatechopper):

ishaan can u give me questions like u used to? post em

OpenStudy (anonymous):

\[\frac{\sqrt{3}}{4}x = \frac{\sqrt{81}}{63}y + \sqrt{121}\] Standard and Slope Intercept

OpenStudy (karatechopper):

no square roots